Danny's father is 4 times as old as Danny. Five years ago, Danny's father was 5 times older. How old is Danny?
Answer by richwmiller(17219)

f=4d,
f-5=5*(d-5)
4d-5=5d-25
20=d
f=80
75=5*15
75=75
ok